In a triangle, the second angle measures four times the first. The measure of the third angle is 18° degrees more than the measure of the second. Find the measures of the three angles.


Found 2 solutions by Alan3354, addingup:
Answer by Alan3354(69443)

In a triangle, the second angle B measures four times the first A. The measure of the third angle C is 18° degrees more than the measure of the second. Find the measures of the three angles.
------------
B = 4A
C = B + 18 = 4A+18
---
A + 4A + 4A+18 = 180 degs
etc

Answer by addingup(3677)

A + B + C = 180
B = 4A
C = B + 18 = 4A + 18
----------------------------------
A + 4A + 4A + 18 = 180
9A = 162
A = 162/9 = 18
B = 4A = 72
C = B + 18 = 72+18 = 90
------------------------------------------
Check:
A + B + C = 180
18 + 72 + 90 = 180 Correct